# Moment Tensor Potential and Equivariant Tensor Network Potential with explicit dispersion interactions

### v1

@@ Abstract
In this study, we investigate the effect of incorporating explicit dispersion interactions in the functional form of machine learning interatomic potentials (MLIPs), particularly in the Moment Tensor Potential and Equivariant Tensor Network potential for accurate modeling of liquid carbon tetrachloride, methane, and toluene. We show that explicit incorporation of dispersion interactions via D2 and D3 corrections significantly improves the accuracy of MLIPs when the cutoff radius is set to a commonly used value of 5—6 Å. We also show that for carbon tetrachloride and methane, a substantial improvement in accuracy can be achieved by extending the cutoff radius to 7.5 Å. However, for accurate modeling of toluene, explicit incorporation of dispersion remains important. Furthermore, we find that MLIPs incorporating dispersion interactions via D2 reach a close level of accuracy to those incorporating D3, and D2 is suitable for accurate modeling of the systems in the study, while being less computationally expensive. We evaluated the accuracy of MLIPs in dimer binding curves compared to ab initio data and in predicting density and radial distribution functions compared to experiments.
@@ I Introduction
Machine learning interatomic potentials (MLIPs) have emerged as widely recognized tools for modeling large systems over an extended timescale with the accuracy of ab initio methods. MLIPs typically achieve relative errors in energy and forces of a few percent, evaluated with respect to ab initio results, while also demonstrating a favorable scaling of computational complexity as the number of atoms increasesZuo et al. (2020); Kozinsky et al. (2023); Zhang et al. (2025). This makes MLIPs an attractive choice for simulating systems that are, on the one hand, computationally prohibitive for conventional ab initio methods and, on the other hand, beyond the accuracy of (semi-)empirical force fields.
@@
Despite significant advances in the development of MLIPs, several key challenges remain. The primary challenge for MLIPs arises from their locality, which is a consequence of the introduction of a cutoff radius that restricts the number of interacting atomsZhang et al. (2025). Although this approximation works well for metals and alloys, it requires greater caution for low-dimensional materials, materials with defects, or moleculesBabaei and Sadeghi (2024); Behler and Csányi (2021); Anstine and Isayev (2023). Moreover, the locality of MLIPs restricts them from accurately capturing long-range interactions, such as electrostatic, induction, and dispersionAnstine and Isayev (2023).
@@
MLIPs that explicitly incorporate long-range interactions have already been presented in the literature. Such MLIPs include PhysNetUnke and Meuwly (2019), TensorMolYao et al. (2018), and ee4G-HDNNP model Ko et al. (2023), which incorporate both electrostatic and dispersion interactions. Other MLIPs such as HDNNP potentialsArtrith, Morawietz, and Behler (2011); Ko et al. (2021), and DPLR modelZhang et al. (2022) incorporate electrostatic interactions. At the same time, GAPBartók et al. (2010) was applied to study systems with strong dispersion interactionsVeit et al. (2019); Rowe et al. (2020, 2018); Ibragimova et al. (2025). However, it is also possible to model intramolecular and intermolecular interactions using separate potentialsWengert et al. (2021); Gao and Remsing (2022). Additionally, an alternative approach is to account for long-range interactions via global descriptorsKabylda et al. (2023).
@@
The most common approach to address long-range interactions is to explicitly incorporate physically motivated terms in the total energy expressionAnstine and Isayev (2023). Behler proposed a classification of MLIPs into four generations based on how long-range interactions are incorporatedBehler (2021); Behler and Csányi (2021). The first and second generations correspond to purely local potentials, while the third generation MLIPs explicitly include long-range interactions, with their strength defined by local features such as atomic charges, dependent on the local environment. The fourth generation accounts for the nonlocality of parameters defining long-range interactions, particularly by making charges dependent on the nonlocal environment. This classification is widely used in the context of MLIPs that incorporate electrostatic interactions and can also be applied to classify MLIPs that account for dispersion interactionsBehler and Csányi (2021).
@@
For explicit capture of dispersion interactions, the most common approach is to utilize one of the dispersion corrections developed for density functional theory (DFT) calculationsGrimme et al. (2016); Hermann, DiStasio Jr, and Tkatchenko (2017), such as Tkatchenko-Scheffler (TS) Tkatchenko and Scheffler (2009), exchange-dipole moment (XDM)Becke and Johnson (2005), D3Grimme et al. (2010) or D2Grimme (2006) models. For electronic density-dependent dispersion correction schemes, such as TS or XDM, machine learning can be used to predict the most computationally demanding parameters, such as atomic moments and volumes in XDMTu et al. (2023, 2024). Further optimization of computational cost can be achieved using averaged parameters instead of machine learning: atomic volumes for TSVeit et al. (2019); Muhli et al. (2021); Ibragimova et al. (2025) or moments and atomic volumes for XDMTu et al. (2023). The latter approach has been shown to have accuracy close to the one of the scheme using machine-learned coefficientsTu et al. (2023). The D2 or D3 dispersion corrections, which do not require electronic density for evaluation, can be applied on top of a local potentialWang and Huang (2020); Ying and Fan (2023); Unke and Meuwly (2019); Yao et al. (2018), without machine learning of the dispersion correction parameters. An alternative involves parameterizing the Lennard-Jones potential based on relevant configurations, such as the exfoliation curve for black phosphorus or grapheneRowe et al. (2018); Deringer, Caro, and Csányi (2020).
@@
In the present work, we focus on the development of MLIPs with explicit treatment of the dispersion interaction, particularly Moment Tensor Potentials (MTPs)Shapeev (2016) and Equivariant Tensor Network (ETN) potentialsHodapp and Shapeev (2024). MTP is implemented in the MLIP-2 software, together with the AL approach based on the D-optimality criterionNovikov et al. (2020), which enables efficient training set generation for MTP. ETN is implemented in the MLIP-4 softwareShapeev (2024). MTP has been widely applied to the modeling of inorganic materialsKorotaev et al. (2019); Podryabinkin et al. (2019); Wang et al. (2020), metals and alloysMarmolejo-Tejada and Mosquera (2022); Tasnadi et al. (2021); Gil and Oda (2022); Zheng et al. (2023), and molten saltsRybin et al. (2024); Sun et al. (2024); Attarian, Morgan, and Szlufarska (2022). However, it was not extensively used to study molecular compoundsRybin, Novikov, and Shapeev (2025); Klimanova, Rybin, and Shapeev (2024); Reicht et al. (2024); Novikov et al. (2024). At the same time, ETN has recently been developedHodapp and Shapeev (2024) and has been successfully applied to defects in metallic alloysHodapp (2025), but has not yet been widely benchmarked in various applications.
Here, our objective is to improve the accuracy and applicability of the MTP and ETN models by explicitly incorporating dispersion interactions via D2 and D3 corrections. In our study, we compare the precision of local MLIPS (MLIPS(l)) with MLIPS with explicit dispersion correction (MLIP+D). We study the accuracy differences between MLIP(l)’s and MLIP+D’s when the cutoff radius is set to a value within a commonly used range of 5—6 Å, thereby constraining the MLIP(l)’s ability to capture the full range of dispersion interactions. We additionally test an alternative approach in which we simply extend the cutoff radius for MLIP(l)’s to 7.5 Å, allowing them to describe the nearly full range of dispersion interactions, and compare the accuracy of the resulting MLIP(l)’s with MLIP+D’s.
This comparative analysis was performed on liquid carbon tetrachloride, methane, and toluene simulations, since these systems exhibit predominant dispersion interactions and are widely represented in the literature. Moreover, carbon tetrachloride and toluene are common nonpolar solvents. To examine the predictive power of potentials, we calculate binding curves and compare them with the ones obtained with ab initio calculations. We also calculate density and radial distribution function (RDF) with the fitted MLIPs and compare them to the experimental ones.
@@ II.2 Equivariant Tensor Network Potentials
$\ldots$.… .
The main difference between ETNs and conventional tensor networks is the implementation of symmetry constraints that render the ETN invariant under the actions of the corresponding symmetry group. In our case of interatomic potentials, we require that the energy per atom $V({\bf\mathfrak{n}}_{i})$ remains invariant under the actions of the group of rotations SO(3). [^1] To encode the full O(3) invariance (rotations and reflections) into ETNs, it suffices that $V({\bf\mathfrak{n}}_{i})$ is a real quantity (more details are provided in the original paperHodapp and Shapeev (2024), section 3.4.2) To that end, we require the feature vectors to be SO(3)-equivariant vectors, that rotate correspondingly with a basis change under SO(3). In the following, we consider a decomposition of each $v$ into an irreducible covariant representation of SO(3) using spherical harmonics. Thus, each index $k$ of a feature vector $v_{k}$ is a multi-index $k=(\ell mn)$, with $\ell=0,\ldots,L$ being the index of the subspace of the irreducible representation, $m\in\{-\ell,-\ell+1,\ldots,\ell\}$ being the dimension of the subspace, and $n=1,\ldots,N(\ell)$ is the number of radial channels corresponding to each $\ell$. We point out that we intentionally deviate from the ordering $n\ell m$, commonly used in quantum physics, that puts the index $n$ in front of $\ell$ and $m$. We have chosen this notation because each multi-index in our tensor network may depend on a different $\ell$. Therefore, since $m$ and $n$ are always assumed to depend on $\ell$, the ordering $\ell mn$ appears to be more comprehensible in our context.
@@
When comparing the moment tensor descriptors $M_{\mu,\nu}({\mathfrak{\bm{n}}}_{i})$ and the feature vectors of ETNs, the radial parameters $c$ are decomposed into $A$ and $B$, and the angular part $r_{ij}^{\otimes\nu}$ is reshaped to $Y_{\ell m}(r_{ij}/|r_{ij}|)$. More details on the relation between MTP and ETN are provided in the original paper by Hodapp and ShapeevHodapp and Shapeev (2024)
@@ II.3 Fitting
For finding parameters of MLIPs, e.g. MTP or ETN parameters, we fit it on a (quantum-mechanical) training set. Let $K$ be a number of configurations in the training set and $N$ be a number of atoms. Denote a vector of MTP/ETN parameters to be found by ${\bm{\theta}}$. We find the optimal parameters ¯${\bm{\bar{}}{\theta}}$ by solving the following optimization problem (minimization of the objective function)
$\begin{array}[]{c}\displaystyle\sum\limits_{k=1}^{K}\Bigl{[}w_{\rm e}\left(E^{\rm MLIP}_{k}({\bm{\theta}})-E^{\rm QM}_{k}\right)^{2}+\\ \displaystyle w_{\rm f}\sum_{i=1}^{N}\left|{\bf f}^{\rm MLIP}_{i,k}({\bm{\theta}})-{\bf f}^{\rm QM}_{i,k}\right|\Bigr{]}^{2}\to\operatorname{min}.\end{array}$ (9)
We start from randomly initialized MLIP parameters. The optimal parameters ¯${\bm{\bar{}}{\theta}}$ are found numerically, using the iterative method to minimize the non-linear objective function, namely the Broyden-Fletcher-Goldfarb-Shanno algorithm. Thus, after optimization, the parameters ¯${\bm{\bar{}}{\theta}}$ are near the local minimum of the objective function. In this objective function, $E^{\rm QM}_{k}$ and ${\bf f}^{\rm QM}_{i,k}$ are the reference energies and forces, i.e., the ones to which we fit the MLIP energies $E^{\rm MLIP}_{k}$ and forces ${\bf f}^{\rm MLIP}_{i,k}$ and thus optimize the MLIP parameters ${\bm{\theta}}$. The factors $w_{\rm e}$ and $w_{\rm f}$ are non-negative weights which express the importance of energies and forces with respect to each other. We took $w_{\rm e}=1$ and $w_{\rm f}=0.01$. We refer to the minimization problem (9) as MLIP fitting.
@@ II.4 Active Learning
Assume that we found the vector of the optimal MTP parameters ¯${\bm{\bar{}}{\theta}}$ after solving (9) and let the length of the vector be $m$. We then construct a matrix
@@ II.5 Range separation
The explicit incorporation of dispersion via D2 or D3 allows to maintain a cutoff radius of 5—6 Å for the machine learning part of these potentials, while the cutoff radius for D2 or D3 is set to 15 Å for all simulations presented in this work.
@@ II.7 D3 correction
where $L_{pq}$ is defined as:
@@ II.8 Protocol of datasets generation and training of MLIPs
Figure 1: The general scheme for generating training sets for all the MLIPs considered. At the first step, the AIMD simulation is conducted and initial MTP is trained on the configurations uniformly chosen from the AIMD trajectory. Next, a pretraining set formed from uniformly selected configurations and configurations, additionally selected from the AIMD trajectories with the MaxVol algorithmPodryabinkin and Shapeev (2016). The initial MTP retrained on the resulting pretraining set. After that the AL of MTP during MD simulations is carried out, and results in the actively selected training set with total energies and forces. Further steps correspond to training of MTP and ETN on the dataset including total energies and forces, and subtracting D2 or D3 contributions and utilizing the resulting training sets for fitting of MTP or ETN part of MTP+D2, ETN+D2, ETN+D3 potentials.
We started creating a training set from ab initio molecular dynamics (AIMD) simulations. An initial configuration was generated using Packmol Martínez et al. (2009). Next, we uniformly chose about 20 configurations from this AIMD trajectory to avoid correlations between configurations and fit an initial MTP. After that, we selected configurations using the MaxVol algorithm from the pool of configurations generated during AIMD, updated the initial training set, and refitted the MTP, obtaining pretrained MTP. Subsequently, we actively trained the pretrained MTP during 10 parallel MD simulations of 100 ps each using the AL procedure described above, utilizing LAMMPS softwareThompson et al. (2022) for MD. As a result, we obtained the actively selected training set with total energies and forces. During both AIMD and MD in the AL procedure, a timestep was set to 1 fs.

### v2

@@ Abstract
In this study, we investigate the effect of incorporating explicit dispersion interactions in the functional form of machine learning interatomic potentials (MLIPs), particularly in the Moment Tensor Potential and Equivariant Tensor Network potential for an accurate modeling of liquid carbon tetrachloride, methane, and toluene. We show that explicit incorporation of dispersion interactions via D2 and D3 corrections significantly improves the accuracy of MLIPs when the cutoff radius is set to the commonly used value of 5–6 Å. We also show that for carbon tetrachloride and methane, a substantial improvement in accuracy can be achieved by extending the cutoff radius to 7.5 Å. However, for accurate modeling of toluene, explicit incorporation of dispersion remains important. Furthermore, we find that MLIPs incorporating dispersion interactions via D2 reach a close level of accuracy to those incorporating D3, implying that D2 is suitable for an accurate modeling of the systems in the study, while being less computationally expensive. We benchmarked the accuracy of the MLIPs on dimer binding curves compared to ab initio data and on predicting density and radial distribution functions compared to experiments.
@@ I Introduction
Machine learning interatomic potentials (MLIPs) have emerged as widely recognized tools for modeling large systems over extended length- and timescales with the accuracy of ab initio methods Zuo et al. (2020); Kozinsky et al. (2023); Zhang et al. (2025). This makes MLIPs an attractive choice for simulating systems that are, on the one hand, computationally prohibitive for conventional ab initio methods and, on the other hand, beyond the accuracy of (semi-)empirical force fields.
@@
Despite significant advances in the development of MLIPs, several key challenges remain. The primary challenge for MLIPs arises from their locality, which is a consequence of the introduction of a cutoff radius that restricts the number of interacting atoms Zhang et al. (2025). Although this approximation works well for metals and alloys, it requires greater caution for low-dimensional materials, materials with defects, or molecules Babaei and Sadeghi (2024); Behler and Csányi (2021); Anstine and Isayev (2023). Moreover, the locality of MLIPs restricts them from accurately capturing long-range interactions, such as electrostatic, induction, and dispersion Anstine and Isayev (2023), interactions.
@@
MLIPs that explicitly incorporate long-range interactions have already been presented in the literature. Such MLIPs include PhysNet Unke and Meuwly (2019), AIMNet2 Isayev, Anstine, and Zubaiuk (2025), SpookyNet Unke et al. (2021, 2024), TensorMol Yao et al. (2018), and ee4G-HDNNP model Ko et al. (2023), which incorporate both, electrostatic and dispersion interactions. Other MLIPs such as HDNNP potentials Artrith, Morawietz, and Behler (2011); Ko et al. (2021), and the DPLR model Zhang et al. (2022) incorporate electrostatic interactions. At the same time, GAP Bartók et al. (2010) was applied to study systems with strong dispersion interactions Veit et al. (2019); Rowe et al. (2020, 2018); Ibragimova et al. (2025). However, it is also possible to model intramolecular and intermolecular interactions using separate potentials Wengert et al. (2021); Gao and Remsing (2022). Additionally, an alternative approach is to account for long-range interactions via global descriptors Kabylda et al. (2023).
@@
The most common approach to address long-range interactions is to explicitly incorporate physically motivated terms in the total energy expression Anstine and Isayev (2023). Behler proposed a classification of MLIPs into four generations based on how long-range interactions are incorporated Behler (2021); Behler and Csányi (2021). The first and second generations correspond to purely local potentials, while the third generation MLIPs explicitly include long-range interactions, with their strength defined by local features such as atomic charges, dependent on the local environment. The fourth generation accounts for the nonlocality of parameters defining long-range interactions, particularly by making charges dependent on the nonlocal environment. This classification is widely used in the context of MLIPs that incorporate electrostatic interactions, and can also be applied to classify MLIPs that account for dispersion interactions Behler and Csányi (2021).
@@
To explicitly capture dispersion interactions, the most common approach is to utilize one of the dispersion corrections developed for density functional theory (DFT) calculations Grimme et al. (2016); Hermann, DiStasio Jr, and Tkatchenko (2017), such as Tkatchenko-Scheffler (TS) Tkatchenko and Scheffler (2009), exchange-dipole moment (XDM) Becke and Johnson (2005), D3 Grimme et al. (2010), or D2 Grimme (2006) models. For electronic density-dependent dispersion correction schemes, such as TS or XDM, machine learning can be used to predict the most computationally demanding parameters, such as atomic moments and volumes in XDM Tu et al. (2023, 2024). Further optimization of the computational cost can be achieved by using averaged parameters instead of machine learning: atomic volumes for TS Veit et al. (2019); Muhli et al. (2021); Ibragimova et al. (2025) or moments and atomic volumes for XDM Tu et al. (2023). The latter approach has been shown to have an accuracy close to that of the scheme using machine-learned coefficients Tu et al. (2023). The D2 or D3 dispersion corrections, which do not require electronic density for evaluation, can be applied on top of a local potential Wang and Huang (2020); Ying and Fan (2023); Unke and Meuwly (2019); Yao et al. (2018), without machine learning of the dispersion correction parameters. An alternative involves parameterizing the Lennard-Jones potential based on relevant configurations, such as the exfoliation curve for black phosphorus or graphene Rowe et al. (2018); Deringer, Caro, and Csányi (2020).
@@
In the present work, we focus on the development of MLIPs with explicit treatment of the dispersion interaction, particularly Moment Tensor Potentials (MTPs) Shapeev (2016) and Equivariant Tensor Network (ETN) potentials Hodapp and Shapeev (2024). MTP is implemented in the MLIP-2 package, together with the AL approach based on the D-optimality criterion Novikov et al. (2020), which enables efficient training set generation for MTP. ETN is implemented in the MLIP-4 package Shapeev (2024). MTP has been widely applied to the modeling of inorganic materials Korotaev et al. (2019); Podryabinkin et al. (2019); Wang et al. (2020), metals and alloys Marmolejo-Tejada and Mosquera (2022); Tasnadi et al. (2021); Gil and Oda (2022); Zheng et al. (2023), and molten salts Rybin et al. (2024); Sun et al. (2024); Attarian, Morgan, and Szlufarska (2022). However, it was not extensively used to study molecular compounds Rybin, Novikov, and Shapeev (2025); Klimanova, Rybin, and Shapeev (2024); Reicht et al. (2024); Novikov et al. (2024). At the same time, ETN has recently been developed Hodapp and Shapeev (2024) and has been successfully applied to defects in metallic alloys Hodapp (2025), but has not yet been widely benchmarked in various applications.
Here, our objective is to improve the accuracy and applicability of the MTP and ETN models by explicitly incorporating dispersion interactions via D2 and D3 corrections. In our study, we compare the precision of local MLIPS (MLIPS(l)) with MLIPS with explicit dispersion correction (MLIP+D). We study the accuracy differences between MLIP(l)’s and MLIP+D’s when the cutoff radius is set to a value within a commonly used range of 5–6 Å, thereby constraining the MLIP(l)’s ability to capture the full range of dispersion interactions. We additionally test an alternative approach in which we simply extend the cutoff radius for MLIP(l)’s to 7.5 Å, allowing them to describe the major part of dispersion interactions, and compare the accuracy of the resulting MLIP(l)’s with MLIP+D’s.
This comparative analysis was performed on liquid carbon tetrachloride, methane, and toluene simulations, since these systems exhibit predominant dispersion interactions and are widely represented in the literature. Moreover, carbon tetrachloride and toluene are common nonpolar solvents. To examine the predictive power of potentials, we calculate the binding curves and compare them with the ones obtained with ab initio calculations. We also calculate the density and radial distribution function (RDF) with the fitted MLIPs and compare them to the experimental ones.
This study is devoted to benchmarking the above MLIPs fitted to PBE+D2/D3 on the density, binding curves, and RDF, as the aim of this study is to investigate the change in accuracy of MLIPs after incorporating explicit dispersion interactions in their functional form. We do not consider more complex properties, such as diffusion coefficients, whose accuracy may be affected by the level of theory or quantum nuclear effects, as this investigation is beyond the scope of this work.
@@ II.2 Equivariant Tensor Network Potentials
$\ldots.$
The main difference between ETNs and conventional tensor networks is the implementation of symmetry constraints that render the ETN invariant under the actions of the corresponding symmetry group. In our case of interatomic potentials, we require that the energy per atom $V({\bf\mathfrak{n}}_{i})$ remains invariant under the actions of the group of rotations SO(3). [^1] To encode the full O(3) invariance (rotations and reflections) into ETNs, it suffices that $V({\bf\mathfrak{n}}_{i})$ is a real quantity (more details are provided in the original paper Hodapp and Shapeev (2024), section 3.4.2) To that end, we require the feature vectors to be SO(3)-equivariant vectors, that rotate correspondingly with a basis change under SO(3). In the following, we consider a decomposition of each $v$ into an irreducible covariant representation of SO(3) using spherical harmonics. Thus, each index $k$ of a feature vector $v_{k}$ is a multi-index $k=(\ell mn)$, with $\ell=0,\ldots,L$ being the index of the subspace of the irreducible representation, $m\in\{-\ell,-\ell+1,\ldots,\ell\}$ being the dimension of the subspace, and $n=1,\ldots,N(\ell)$ is the number of radial channels corresponding to each $\ell$. We point out that we intentionally deviate from the ordering $n\ell m$, commonly used in quantum physics, that puts the index $n$ in front of $\ell$ and $m$. We chose this notation because each multi-index in our tensor network may depend on a different $\ell$. Therefore, since $m$ and $n$ are always assumed to depend on $\ell$, the ordering $\ell mn$ appears to be more comprehensible in our context.
@@
When comparing the moment tensor descriptors $M_{\mu,\nu}({\mathfrak{\bm{n}}}_{i})$ and the feature vectors of ETNs, the radial parameters $c$ are decomposed into $A$ and $B$, and the angular part $r_{ij}^{\otimes\nu}$ is reshaped to $Y_{\ell m}(r_{ij}/|r_{ij}|)$. More details on the relation between MTP and ETN are provided in the original paper by Hodapp and Shapeev Hodapp and Shapeev (2024).
@@ II.3 Fitting
For finding parameters of MLIPs, e.g. MTP or ETN parameters, we fit it on a (quantum-mechanical) training set. Let $K$ be a number of configurations in the training set and $N$ be a number of atoms. Denote a vector of MTP/ETN parameters to be found by $\bm{\theta}$}. We find the optimal parameters $\bm{\bar{}}{\theta}$} by solving the following optimization problem (minimization of the objective function)
$\begin{array}[]{c}\displaystyle\sum\limits_{k=1}^{K}\Bigl{[}w_{\rm e}\left(E^{\rm MLIP}_{k}({\bm{\theta}})-E^{\rm QM}_{k}\right)^{2}+\\ \displaystyle w_{\rm f}\sum_{i=1}^{N}\left|{\bf f}^{\rm MLIP}_{i,k}({\bm{\theta}})-{\bf f}^{\rm QM}_{i,k}\right|{{}^{2}}\Bigr{]}\to\operatorname{min}.\end{array}$ (9)
We start from randomly initialized MLIP parameters. The optimal parameters $\bm{\bar{}}{\theta}$} are found numerically, using the iterative method to minimize the non-linear objective function, namely the Broyden-Fletcher-Goldfarb-Shanno algorithm. Thus, after optimization, the parameters $\bm{\bar{}}{\theta}$} are near the local minimum of the objective function. In this objective function, $E^{\rm QM}_{k}$ and ${\bf f}^{\rm QM}_{i,k}$ are the reference energies and forces, i.e., the ones to which we fit the MLIP energies $E^{\rm MLIP}_{k}$ and forces ${\bf f}^{\rm MLIP}_{i,k}$ and thus optimize the MLIP parameters $\bm{\theta}$}. The factors $w_{\rm e}$ and $w_{\rm f}$ are non-negative weights which express the importance of energies and forces with respect to each other. We took $w_{\rm e}=1$ and $w_{\rm f}=0.01$ to fit the ETN-based models and we took $w_{\rm e}=1/N$ and $w_{\rm f}=0.01$ to train the MTP-based models. These weights are default in the MLIP-2 and MLIP-4 codes, respectively. We refer to the minimization problem (9) as MLIP fitting.
@@ II.4 Active Learning
Assume that we found the vector of the optimal MTP parameters $\bm{\bar{}}{\theta}$} after solving (9) and let the length of the vector be $m$. We then construct a matrix
@@ II.5 Range separation
The explicit incorporation of dispersion via D2 or D3 allows to maintain a cutoff radius of 5–6 Å for the machine learning part of these potentials, while the cutoff radius for D2 or D3 is set to 15 Å for all simulations presented in this work.
@@ II.7 D3 correction
It is possible to include the three-body interaction term in the expression of the D3 energy. However, in our implementation, we limited D3 to the two-body form, as this implementation is widely used and the contribution of the three-body interaction term is typically much lower than the two-body term Grimme et al. (2016).
@@
where $N_{i}$, $N_{j}$ are numbers of reference compounds for elements $i$ and $j$, and $L_{pq}$ is defined as:
@@ II.8 Protocol of datasets generation and training of MLIPs
Figure 1: The general scheme for generating training sets for all the MLIPs considered. At the first step, the AIMD simulation is conducted and initial MTP is trained on the configurations uniformly chosen from the AIMD trajectory. Next, a pretraining set formed from uniformly selected configurations and configurations, additionally selected from the AIMD trajectories with the MaxVol algorithm Podryabinkin and Shapeev (2016). The initial MTP retrained on the resulting pretraining set. After that, the AL of MTP during MD simulations is carried out, and results in the actively selected training set with total energies and forces. Further steps correspond to training of MTP and ETN on the dataset including total energies and forces, and subtracting D2 or D3 contributions and utilizing the resulting training sets for fitting of MTP or ETN part of MTP+D2, ETN+D2, ETN+D3 potentials.
We started creating a training set from ab initio molecular dynamics (AIMD) simulations. An initial configuration was generated using Packmol Mart´ınez et al. (2009). Next, we uniformly chose about 20 configurations from this AIMD trajectory to avoid correlations between configurations and fit an initial MTP. After that, we selected configurations using the MaxVol algorithm from the pool of configurations generated during AIMD, updated the initial training set, and refitted the MTP, obtaining pretrained MTP. Subsequently, we actively trained the pretrained MTP during 10 parallel MD simulations (AL-MD) of 100 ps each using the AL procedure described above, utilizing LAMMPS package Thompson et al. (2022) for MD. As a result, we obtained the actively selected training set with total energies and forces. During both AIMD and MD in the AL procedure, a timestep was set to 1 fs.